# Photonic quantum generative adversarial networks for classical data

### v1

@@ Abstract
When Generative Adversarial Networks (GANs) first emerged, they marked a breakthrough in the field of classical machine learning. Researchers have since designed quantum versions of the algorithm, both for the generation of classical and quantum data, but most work so far has focused on qubit-based architectures. In this article, we focus on photonic quantum computing and present a quantum GAN based on linear optical circuits and Fock-space encoding for the generation of classical data. We explore the trainability and the performance of the model in a proof-of-concept image generation scenario. We then conduct an experiment where we train our quantum GAN on Quandela’s photonic quantum processor Ascella.
@@ I Introduction
Our work sets out to explore photonic quantum computing for the generation of classical data. In particular, we propose a quantum generative adversarial network (QGAN) where the generator network is a variational quantum circuit [12]. We train our model on the MNIST [13] dataset of handwritten digits in reduced dimension, using a patch-based image generation approach, in both ideal and noisy settings. Additionally, we run the full training procedure as a physical experiment on Quandela’s quantum processing unit Ascella introduced in [9]. Our work is a proof-of-concept demonstration that photonic quantum adversarial models can be trained to generate classical data.
@@ II.1 Classical GANs
In practice, a batch of noise samples from the latent space is supplied to the generator. It produces a batch of results, which are then used for the discriminator. Simultaneously, the discriminator is supplied a batch of samples from the training dataset. The loss is therefore computed by averaging over the batch, which has the added advantage of stabilizing the learning process. For every learning iteration, optimization steps are performed first for the discriminator, and then for the generator, using gradient descent (or ascent), where the gradient is computed using backpropagation. The number of optimization steps $k$ for the discriminator depends on the specific use case. The full training algorithm is shown Appendix A.
@@ III.3 Training and optimization
The training of photonic QGANs progresses similarly to the regular GANs, as in Algorithm 1. The classical discriminator is trained first for one step using backpropagation-enabled stochastic gradient ascent to maximize $L_{D}$, after which the quantum generator is trained for several steps of Simultaneous Perturbation Stochastic Approximation (SPSA) [35] iterations, to optimize $L_{G}$. All the parameters are updated, and this process is repeated until the maximum number of training epochs is reached.
@@
III.4 Numerical experiments: ideal simulations
We perform our numerical experiments using Quandela’s software package Perceval [38], designed for linear optical quantum circuits. We include an abstract pseudocode in Appendix A, and our Python code can be found in a companion Github repository [39].
Ideal conditions for simulations assume a perfect single photon source, ideal components, no photon loss, perfect detectors and the absence of sampling errors. First, we focus on the generation of digit "0" for the design of the model and the optimization of its hyperparameters. We then further study the best model for the generation of other digits, and for noisy simulations in the next sections. We include all details about the model optimization and hyperparameter search in Appendix B.
@@ III.4 Numerical experiments: ideal simulations
It appears that there is no specific loss value where the training can be stopped. However, good models generally reach an equilibrium where $L_{G}$ and $L_{D}$ losses start to fluctuate around the same moving average, without increasing or decreasing further. This is clearly observed in Figure 6c where both configurations converge to an equilibrium after around 1000 iterations. This equilibrium is centered around value close to $0.69$, which corresponds to $\log 2$. Taking a look at loss from Equation (4), it becomes clear that $L_{G}=-\log 2$ , when $D(G(z))=1/2$. That is, when the discriminator prediction about whether the fake image is real is as good as a random guess. Despite the fact this equilibrium is reached around iteration 1000, we note that further training does improve the results. Therefore, cutting off the training once the losses reach the value of $\log 2$ is not necessarily a desirable strategy. Rather, it makes sense to inspect the generated results using qualitative or quantitative metrics in order to be able to cutoff the training in due time. Quantitative assessment metrics are not discussed here, since the small scale of generated images allows for direct inspection of the results, as has been the case in many quantum generative learning works [26].
We refer the reader to Appendix B for further analysis of the results and for hypotheses as to why some configurations perform better than others. Overall, the models from Figure 6c behaved best and we select them for further simulations.
@@
In particular, we apply the model from the left hand side of Figure 6c to the generation of other digits. We present results for some digits in Figure 7, while the rest of the digits can be found in the Github respository [39]. An additional 500 iterations were employed here in order to properly assess the convergence for different digits. When comparing to target data, we see that the model performs fairly well for most of the digits, and each sampled digit has recognizable contours. Importantly, the model either converges or is bound to do so for all the digits, which is indicated by a narrowing of the standard deviation for the $L_{G}$ loss. While some digits may require more training iterations to reach an equilibrium state, we can assert that the model starts producing realistic results around iteration 1500.
@@ III.5 Noisy simulations
Clearly, the added noise considerably slows down the convergence. While the noiseless version reaches an equilibrium around iteration 900 on average, for the noisy version on the right of Figure 8, even 1500 iterations are not enough. This version, which has more sub-generators, shows a diverging learning trend. This might be caused by the increased number of parameters, making it harder to train this model under the predefined learning rate restrictions (see Appendix B). Nevertheless, for the model on the left which has less sub-generators but a higher number of modes, one can see how the standard deviation of the generator decreases throughout the training, indicating that it is likely to eventually converge. Moreover, the generated results are satisfactory when applying a manual check. Results comprehensible to the human eye are available by iteration 600 for the ideal version, while the noisy version requires more than 1000.
@@ III.6 Physical experiment
Based on the insights gathered from our simulations, we run an experiment on Quandela’s processor Ascella, with our best performing model. The corresponding quantum circuit is of a reasonable depth, making it compatible with Ascella’s chip, which is naturally reused for each sub-generator. In order to make the duration of the experiment tractable, we decrease the total number of training iterations to 1000, with 3 SPSA steps for the generator at each iteration, which results in a total of 3000 SPSA steps. The results for one training instance are displayed in Figure 9.
@@ IV Discussion and future work
Most of the QGAN literature concentrates on smaller scale models for generation of lower resolution images, such as $3\times 3$ bars-and-stripes. However, the task of generating larger images is tractable with current hardware if the right approach is used – such as our distribution-based mapping. This highlights the importance of looking for alternative strategies when working with near-term quantum hardware. Using our mapping along with noise reuploading gave us greater flexibility, and in combination with patch-based learning the model could generate higher resolutions images.
We note that SPSA-based optimization seems to work surprisingly well on variational linear optical circuits. Especially in the noisy experimental setup, SPSA allowed for stable learning at a constant cost, albeit requiring more iterations to converge. Fine-tuning the optimization and the model in general was an important aspect of the work. Future work may concentrate on improving the convergence rates of optimization techniques for photonic platforms, which can be used directly to improve the performance of our model. Indeed, only few works so far have focused on exploring methods like parameter-shift rules [41] in the context of quantum photonics [42]. As mentioned in the text, other improvements to this work include the extension of the model to a conditional QGAN, as well as the integration of error mitigation techniques for sources of noise such as photon loss and distinguishability [43].
@@
As is often the case in quantum machine learning, it is not always clear if a quantum model will perform better than a classical one and on which datasets that would be the case [44], although interesting alternative benchmarks such as the energy consumption of quantum versus classical computers are being explored [45]. Tasks based on quantum data clearly require a quantum model – for example if the generator learns how to implement a quantum gate or generate specific quantum states. For classical data on the other hand, a deeper investigation is required, likely related to the identification of inductive biases of quantum models. In our case in particular, identifying inductive biases of photonic quantum circuits may help us understand where photonic QGANs may be most useful.
@@ Acknowledgements
We would like to thank Pierre-Emmanuel Emeriau for fruitful discussions and feedback on the manuscript. We acknowledge funding from the Plan France 2030 through the project OQuLus ANR-23-PETQ-0013.
@@ Appendix B Model details
In section III.4, we optimize the model in ideal conditions before proceeding with noisy simulations and the experiment. This optimization is done through hyperparameter search, by testing and choosing the models that showed the best potential and performance. Batch size is one of the fixed hyperparameters over our search. All generators are supplied $4$ noise samples and therefore produce a batch of $4$ results over which averaging is done for loss evaluation. This batch size was found to be optimal in terms of computation time and implications on the stability of the training. Another training hyperparameter which is fixed across all models is the learning rate, for both the generator and the discriminator. For the discriminator, we chose a constant learning rate of $0.002$. For the generator, the adaptive learning rate of SPSA is employed, with the initial rate dependent on the initial gradient values. However, the number of optimization steps per each discriminator step is fixed to be $7$. A total of $10500$ SPSA iterations are used, with $7$ steps corresponding to a single step of gradient ascent of the discriminator. This means that the total optimization routine has $10500/7=1500$ steps.
Four hyperparameters were picked to be variable: the number of sub-generators, their circuit structure as shown in Figure 5, their input state, and the PNR capabilities of the detectors. We do a grid search on these configurations, and we present the most interesting results in Figures 6a - 6d from the main text. The full list of the results is presented in the companion Github repository [39] for curious readers.

### v2

@@ Abstract
In generative learning, models are trained to produce new samples that follow the distribution of the target data. These models were historically difficult to train, until proposals such as Generative Adversarial Networks (GANs) emerged, where a generative and a discriminative model compete against each other in a minimax game. Quantum versions of the algorithm were since designed, both for the generation of classical and quantum data. While most work so far has focused on qubit-based architectures, in this article we present a quantum GAN based on linear optical circuits and Fock-space encoding, which makes it compatible with near-term photonic quantum computing. We demonstrate that the model can learn to generate images by training the model end-to-end experimentally on a single-photon quantum processor.
@@ I Introduction
In this work, we propose a quantum generative adversarial network (QGAN) where the generator network is a variational photonic quantum circuit [12]. We train our model on the MNIST [13] dataset of handwritten digits in reduced dimension, using a patch-based image generation approach, in both ideal and noisy settings. We run the full training procedure as a physical experiment on Quandela’s quantum processing unit Ascella introduced in [9], whose setup consists of a single-photon source connected to a integrated photonic chip and photon detectors. Our work is a proof-of-concept demonstration that photonic quantum adversarial models can be trained to generate classical data, and our results thus contribute to a growing body of literature on near-term quantum machine learning implementations.
@@ II.1 Classical GANs
In practice, a batch of noise samples from the latent space is supplied to the generator. It produces a batch of results, which are then used for the discriminator. Simultaneously, the discriminator is supplied a batch of samples from the training dataset. The loss is therefore computed by averaging over the batch, which has the added advantage of stabilizing the learning process. For every learning iteration, optimization steps are performed first for the discriminator, and then for the generator, using gradient descent (or ascent), where the gradient is computed using backpropagation. The number of optimization steps $k$ for the discriminator depends on the specific use case. The full training algorithm is shown in the Supplemental Document in Section S1.
@@ III.3 Training and optimization
The training of photonic QGANs progresses similarly to the regular GANs, as in Algorithm S1. The classical discriminator is trained first for one step using backpropagation-enabled stochastic gradient ascent to maximize $L_{D}$, after which the quantum generator is trained for several steps of Simultaneous Perturbation Stochastic Approximation (SPSA) [35] iterations, to optimize $L_{G}$. All the parameters are updated, and this process is repeated until the maximum number of training epochs is reached.
@@
III.4 Numerical experiments: ideal simulations and assessment
We perform our numerical experiments using Quandela’s software package Perceval [38], designed for linear optical quantum circuits. We include an abstract pseudocode in Supplemental Document in Section S1, and our Python code can be found in a companion Github repository [39].
Ideal conditions for simulations assume a perfect single photon source, ideal components, no photon loss, perfect detectors and the absence of sampling errors. First, we focus on the generation of digit "0" for the design of the model and the optimization of its hyperparameters. We then further study the best model for the generation of other digits, and for noisy simulations in the next sections. We include all details about the model optimization and hyperparameter search in the Supplemental Document in Section S2.
@@ III.4 Numerical experiments: ideal simulations and assessment
It appears that there is no specific loss value where the training can be stopped. However, good models generally reach an equilibrium where $L_{G}$ and $L_{D}$ losses start to fluctuate around the same moving average, without increasing or decreasing further. This is clearly observed in Figure 6c where both configurations converge to an equilibrium after around 1000 iterations. This equilibrium is centered around value close to $0.69$, which corresponds to $\log 2$. Taking a look at loss from Equation (4), it becomes clear that $L_{G}=-\log 2$ , when $D(G(z))=1/2$. That is, when the discriminator prediction about whether the fake image is real is as good as a random guess. Despite the fact this equilibrium is reached around iteration 1000, we note that further training does improve the results. Therefore, cutting off the training once the losses reach the value of $\log 2$ is not necessarily a desirable strategy.
Loss does not constitute a good assessment metric, thus it makes sense to inspect the generated results using other qualitative or quantitative metrics. Important quantitative metrics in the assessment of generative models are the quality and the diversity of generated images. The quality measures the similarity to the original training data, while the diversity measures the variability between the generated images. Essentially, these metrics allow to evaluate a model’s capability to learn well from data instead of memorizing it.
A common choice of measures of similarity and diversity are the Inception Score (IS) [40] and Fréchet Inception Distance (FID) [41], which measure both the quality and the diversity of generated images. However, for a simple model working with lower dimensional images, such as ours, a simpler metric suffices. For this purpose, we chose the Structural Similarity Index Measure (SSIM) [42], which works by weighing in the comparative properties of luminance ($l$), contrast ($c$) and structure ($s$), defined as follows for patches $x$ and $y$ of images to be compared:
$l(x,y)=\frac{2\mu_{x}\mu_{y}+c_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+c_{1}},\quad c(x,y)=\frac{2\sigma_{x}\sigma_{y}+c_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}},\quad$
$s(x,y)=\frac{\sigma_{xy}+c_{3}}{\sigma_{x}\sigma_{y}+c_{3}},$
where $\mu$ and $\sigma$ are correspondingly the mean and the standard deviation of pixel intensity for the given patch, $\sigma_{xy}$ is the covariance between patches $x$ and $y$, and $c_{1}$, $c_{2}$ and $c_{3}$ are stabilizing constants, preventing a division by a small value. With these notions in mind SSIM is defined as:
$\operatorname{SSIM}(x,y)$ $=l(x,y)\cdot c(x,y)\cdot s(x,y)$
$=\frac{\left(2\mu_{x}\mu_{y}+c_{1}\right)\left(2\sigma_{xy}+c_{2}\right)}{\left(\mu_{x}^{2}+\mu_{y}^{2}+c_{1}\right)\left(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{3}\right)}.$
SSIM compares large images patch by patch and then averages the metrics across the patches, however for our case due to the small sizes of images, it compares them immediately. SSIM varies in the range $[-1,1]$, with values close to $1$ indicating similarity, 00 indicating absence of similarity and values close to $-1$ indicating anticorrelation between images.
In order to evaluate the models we compute the similarity score, which computes a pairwise SSIM of 500 random real images from the training dataset against 500 images generated by the quantum generator and then average across all the images. We also compute the the diversity score: first we compute the pairwise SSIM among the generated images, and again average across all images. This value is then subtracted from 1 to follow the logic of high diversity corresponding to more diverse (rather than more similar) images.
To have a baseline for comparison, we train a small GAN with a classical generator, which has approximately the same amount of parameters as the largest generator considered in this work (around 330). This model is one of the simplest possible generators with only 1 small hidden layer, built in a way to allow for a fair comparison with quantum analog, with the training hyperparameters identical to those utilized for quantum generators. We note that a classical generator like this is heavily underparametrized and would not normally be used in a setting other than a comparative one, however it mimics well the behaviour of small quantum models discussed here.
Figure 7: Similarity plotted against diversity of generated images.
We plot the similarity against the diversity for all the models tested, along with corresponding scores of the classical model on a scatter plot presented on Figure 7. Firstly, we note that diversity and similarity are anticorrelated for small models discussed here. In addition, the quantum model with the highest similarity (highlighted as a square, corresponding to Figure 6c, left) has a performance comparable to that of the classical model (highlighted as a triangle), with slightly worse similarity and a slightly better diversity. Moreover, the quantum generator has around $2/3$ parameters of the classical one (208 vs 332).
Visual inspection confirms that indeed the results are of generally high quality for this model, albeit not very diverse. We thus choose this model for further testing, since it compares the best to the classical alternative. We refer the reader to Section S2 of the Supplemental Document for further analysis of the results and for hypotheses as to why some configurations perform better than others.
@@
We apply the model from the left hand side of Figure 6c to the generation of other digits. We present results for some digits in Figure 8, while the rest of the digits can be found in the Github respository [39]. An additional 500 iterations were employed here in order to properly assess the convergence for different digits. When comparing to target data, we see that the model performs fairly well for most of the digits, and each sampled digit has recognizable contours. Importantly, the model either converges or is bound to do so for all the digits, which is indicated by a narrowing of the standard deviation for the $L_{G}$ loss. While some digits may require more training iterations to reach an equilibrium state, we can assert that the model starts producing realistic results around iteration 1500.
@@ III.5 Noisy simulations
Clearly, the added noise considerably slows down the convergence. While the noiseless version reaches an equilibrium around iteration 900 on average, for the noisy version on the right of Figure 9, even 1500 iterations are not enough. This version, which has more sub-generators, shows a diverging learning trend. This might be caused by the increased number of parameters, making it harder to train this model under the predefined learning rate restrictions (see Supplemental Document, Section S2). Nevertheless, for the model on the left which has less sub-generators but a higher number of modes, one can see how the standard deviation of the generator decreases throughout the training, indicating that it is likely to eventually converge. Moreover, the generated results are satisfactory when applying a manual check. Results comprehensible to the human eye are available by iteration 600 for the ideal version, while the noisy version requires more than 1000.
@@ III.6 Physical experiment
Based on the insights gathered from our simulations, we run an experiment on Quandela’s processor Ascella, with our best performing model. The experimental setup is the following: a gated InGaAs quantum dot placed inside a micropillar cavity [44] is excited by a laser as in the scheme of [45] to produce single photons on demand. A demultiplexer converts the train of produced photons into single photons arriving simultaneously on the chip, which is a 12-mode interferometer designed according to the Clements scheme [46]. The phase shifters on the chip are tuned thermo-optically in order to apply the intended operations, and this process is optimised following the characterization procedure of [47]. The photons then enter superconducting nanowire single-photon detectors (SNSPD) detectors, which are threshold detectors. As mentioned in the previous section, the setup is affected by photon loss and distinguishability.
The circuit that corresponds to our best performing model is of reasonable depth, making it compatible with the Ascella chip, which is naturally reused for each sub-generator. In order to make the duration of the experiment tractable, we decrease the total number of training iterations to 1000, with 3 SPSA steps for the generator at each iteration, which results in a total of 3000 SPSA steps. The results for one training instance are displayed in Figure 10.
@@ IV Discussion and future work
Most of the QGAN literature concentrates on smaller scale models for generation of lower resolution images, such as $3\times 3$ bars-and-stripes. However, the task of generating larger images is tractable with currently available hardware if the right approach is used – such as our distribution-based mapping. This highlights the importance of looking for alternative strategies when working with near-term quantum hardware. Using our mapping along with noise reuploading gave us greater flexibility, and in combination with patch-based learning the model could generate higher resolutions images.
In terms of scaling, we point out that our approach requires the approximation of the output probability distribution and may thus have a high sampling cost, like many quantum machine learning models. Photonics at least have the benefit of providing fast sampling which improves runtime, a strength which will become more apparent as levels of photon loss decrease with the advancement of the technology. However, beyond a certain dimension of the problem, it might be necessary to modify the approach, for instance using our model together with a VAE that transforms the data.
We note that, while the size of the Fock space grows exponentially with the number of modes and photons, it is also directly linked to the dimension of the target data, given our mapping. Many problems of interest might thus be reachable with a reasonably sized device. When comparing the native photonic approach to the qubit-based approach of [26], we see how in terms of resources utilised, on photonic platforms it may be advantageous to use the Fock space over the Hilbert space. As we saw above, a model can be built which uses 5 modes and 2 photons for the generation of 8x8 images, with the input state |$\ket{0,1,0,1,0}$, as compared to 8 modes and 4 photons, which would be needed in dual-rail encoding to have qubits as in [26].
Concerning the optimization, we note that SPSA seems to work surprisingly well on variational linear optical circuits. Especially in the noisy experimental setup, SPSA allowed for stable learning at a constant cost, albeit requiring more iterations to converge. Fine-tuning the optimization and the model in general was an important aspect of the work. Future work may concentrate on improving the convergence rates of optimization techniques for photonic platforms, which can be used directly to improve the performance of our model. Indeed, only few works so far have focused on exploring methods like parameter-shift rules [48] in the context of quantum photonics [49, 50, 51]. As mentioned in the text, other improvements to this work include the extension of the model to a conditional QGAN, as well as the integration of error mitigation techniques for sources of noise such as photon loss and distinguishability [52].
@@
In terms of applications, it is not always clear for which problems a quantum model will perform best – a recent study showed that several popular quantum classifiers do not outperform their classical counterparts on standard datasets [53]. A solution that is often proposed is to focus on quantum data instead, e.g., to generate quantum states. However, problems containing classical data are overall much more common. There, the concept of inductive bias proves very helpful to follow, requiring that the model contain some information about the structure of the problem at hand. For this reason, high-energy physics and quantum chemistry are often cited as areas where quantum models might be most useful on classical data. For example, in [54], a QGAN is trained to detect anomalous events that cannot be described by the standard model; in [55] a QGAN was used to generate calorimeter shower images; and in [56, 57], QGANs and QCBMs were used for the generation of new molecules.
We saw that our best performing generator achieves similar performance to its classical counterpart for a lower number of trainable parameters. Such benchmarks can be interesting to focus on, as an alternative to speedups. In the same spirit, the energy consumption of quantum versus classical computers are currently being explored [58].
@@ Acknowledgements
We would like to thank Pierre-Emmanuel Emeriau for fruitful discussions and feedback on the manuscript. This work has been co-funded by the European Commission as part of the EIC program under the grant agreement 101130384 for the Quondensate project. We also acknowledge funding from the Plan France 2030 through the project OQuLus ANR-22-PETQ-0013.
@@ Appendix B Model details
In section III.3 of the main text, we optimize the model in ideal conditions before proceeding with noisy simulations and the experiment. This optimization is done through hyperparameter search, by testing and choosing the models that showed the best potential and performance. Batch size is one of the fixed hyperparameters over our search. All generators are supplied $4$ noise samples and therefore produce a batch of $4$ results over which averaging is done for loss evaluation. This batch size was found to be optimal in terms of computation time and implications on the stability of the training. Another training hyperparameter which is fixed across all models is the learning rate, for both the generator and the discriminator. For the discriminator, we chose a constant learning rate of $0.002$. For the generator, the adaptive learning rate of SPSA is employed, with the initial rate dependent on the initial gradient values. However, the number of optimization steps per each discriminator step is fixed to be $7$. A total of $10500$ SPSA iterations are used, with $7$ steps corresponding to a single step of gradient ascent of the discriminator. This means that the total optimization routine has $10500/7=1500$ steps.
Much like the baseline classical generator discussed in the main body in the assessment of the models, the discriminator is a simple fully connected network, but it has more layers and parameters. With 3 linear layers interleaved with ReLU activation functions and around 5000 parameters initialized randomly, it motivates the faster learning of the generator, by being more accurate in the classification of fake and real images. Since this is a classical model the optimisation is performed using stochastic gradient descent.
Four hyperparameters were picked to be variable: the number of sub-generators, their circuit structure as shown in Figure 5 of the main text, their input state, and the PNR capabilities of the detectors. We do a grid search on these configurations, and we present the most interesting results in Figure 6 from the main text. The full list of the results is presented in the companion Github repository (https://github.com/Quandela/photonic-qgan) for curious readers.